# Learning to Extract Rational Evidence via Reinforcement Learning for Retrieval-Augmented Generation

### v2

@@ 1 Introduction
Recently, several studies have attempted to address this issue. They can be mainly divided into two categories: (1) Reranking more relevant passages go up to the top of the retrieval list Hwang et al. (2024b); Mortaheb et al. (2025); Mao et al. (2021); and (2) Summarizing retrieval contents into a coherent narrative with higher signal to noise ratio Wang et al. (2023); Zhao et al. (2024a); Zhu et al. (2024); Xu et al. (2024). The former heavily relies on the performance of chosen rerankers and may disrupt context after reranking. It usually performs worse than the latter due to a lack of contextual understanding. The latter aims to train LLMs as a filtering model through fine-tuning Wang et al. (2023) or preference optimization Zhao et al. (2024a). They typically create the training data using hand-crafted strategies, e.g., String Inclusion that measures whether the golden answer is included in the candidate passage, and Lexical Overlap, calculating the unigram overlap between the golden answer and candidate passage.
@@ 2.1 Problem Statement
In RAG Lewis et al. (2020); Wu et al. (2022), LLMs are given an input query $q$ and a set of top-$k$ retrieved passages $P=\{p_{1},p_{2},...,p_{k}\}$, which aim to generate an output $o$ closing to the golden answer $a$, conditioned on ($q$, $P$). In the traditional RAG paradigm, top-$k$ retrieved passages are directly fed into LLMs. However, retrieved passages $P$ often contain irrelevant or noisy contents Wang et al. (2023); Zhao et al. (2024a), which considerably degrade LLMs’ generation quality. Given that, an additional evidence extraction model is proposed to condense $P$ into a concise, query-relevant context $e$ to improve generation quality and speed. Formally, the paradigm of RAG with evidence extraction can be formulated as follows:
@@
where $\mathcal{U}(\cdot)$ is a utility function (e.g., $\mathrm{F}_{1}$ score) and is used to measure the quality of outputs conditioned on the golden answer $a$; $|e|$ and $|P|$ denote the length of evidence  and  passages, respectively.
@@ 3.2 Modeling Verifiable Reward
where $R^{ans}$ is the average answer reward of $R^{ans}_{r}$, $R^{ans}_{e}$ and $R^{ans}_{f}$; $R^{fmt}$ is the average length one of $R^{fmt}_{r}$ and $R^{fmt}_{e}$; $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$ are hyperparameters; $R^{final}$ is the final reward used in RL training.
@@ 3.3 Online Policy Optimization
where $\mathrm{mean}(\cdot)$ and $\mathrm{std}(\cdot)$ compute the average and standard deviation, respectively. However, GRPO’s group normalization may overly magnify the minor numerical fluctuation. Taking =$\mathcal{R}=\{0.49,0.51\}$ for example, the $\mathrm{mean}(R)$ is $0.5$, the $\mathrm{std}(R)$ is $0.01$, and the computed advantages are $\{-1.0,1.0\}$, which overly magnifies the minor numerical fluctuation. To mitigate this issue, we propose to clip $\mathrm{std}(R)$ to be at least $\epsilon_{std}$ to ensure that the denominator does not become too small:
@@ 4.1 Experimental Setup
Implementation Details.   We employ Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [^5] https://github.com/QwenLM/Qwen2.5 Yang et al. (2025) as the initial models. We train these two models to extract rational evidence through full parameter fine-tuning for 1 epoch on the NQ dataset. The learning rate is $1e^{-6}$, and we select the best-performing model on the dev set. For modeling verifiable rewards, we tune $\tau$ and $\gamma$ within the ranges of $\{0.1,0.2,0.5,1.0\}$ and $\{0.1,0.3,0.5,0.8,1.0\}$, respectively. For the length threshold $\omega$ and the weighting coefficients (i.e., $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$), we empirically set $\omega$, $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$ as 0.9, 0.8, 0.1, 0.1, respectively, which commonly leads to satisfactory performance. For policy optimization, we set $\epsilon_{std}$ as 0.1 to stabilize training. During training, we set the PPO clipping hyperparameter $\epsilon$ and the coefficient $\beta$ controlling KL-divergence as 0.2 and $1e^{-2}$, respectively. For QA generators, we employ LoRA Yang et al. (2025) to fine-tune Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct to predict the golden answer $a$ based on the query and retrieved passages ($q$, $P$) or only the query (used for the baseline ‘Zero’), where LoRA rank is set as 8.
@@ 5.2 RAG with Context Compression
RAG systems usually concatenate all retrieved passages as the context of LLMs. However, this may introduce data noise and computational overhead due to imperfect retrieval and overlong context. Recently, many works have attempted to compress context and retain important information, mainly including two categories: (1) Reranking methods, which rerank retrieved passages and retain top-ranked passages Hwang et al. (2024b); Mortaheb et al. (2025); Mao et al. (2021) (2) Summarization methods, which extract relevant information from retrieved passages and consolidate them into a narrativeWang et al. (2023); Zhao et al. (2024a); Zhu et al. (2024); Xu et al. (2024). Despite effectiveness, existing works neglect explicit evidence reasoning, which risks filtering out key clues as well as struggles with generalization.

### v3

@@ 1 Introduction
Recently, several studies have attempted to address this issue. They can be mainly divided into two categories: (1) Reranking more relevant passages go up to the top of the retrieval list Hwang et al. (2024b); Mortaheb et al. (2025); Mao et al. (2021); Chen et al. (2025); and (2) Summarizing retrieval contents into a coherent narrative with higher signal to noise ratio Wang et al. (2023); Zhao et al. (2024a); Zhu et al. (2024); Xu et al. (2024). The former heavily relies on the performance of chosen rerankers and may disrupt context after reranking. It usually performs worse than the latter due to a lack of contextual understanding. The latter aims to train LLMs as a filtering model through fine-tuning Wang et al. (2023) or preference optimization Zhao et al. (2024a). They typically create the training data using hand-crafted strategies, e.g., String Inclusion that measures whether the golden answer is included in the candidate passage, and Lexical Overlap that calculates the unigram overlap between the golden answer and candidate passage.
@@ 2.1 Problem Statement
[…]
@@
[…]
@@ 3.2 Modeling Verifiable Reward
where $R^{ans}$ is the average answer reward of $R^{ans}_{r}$, $R^{ans}_{e}$ and $R^{ans}_{f}$; $R^{len}$ is the average length one of $R^{len}_{r}$ and $R^{len}_{e}$; $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$ are hyperparameters; $R^{final}$ is the final reward used in RL training.
@@ 3.3 Online Policy Optimization
where $\mathrm{mean}(\cdot)$ and $\mathrm{std}(\cdot)$ compute the average and standard deviation, respectively. However, GRPO’s group normalization may overly magnify the minor numerical fluctuation. Taking =$\mathcal{R}=\{0.49,0.51\}$ for example, the $\mathrm{mean}(R)$ is 0.50.50.5, the $\mathrm{std}(R)$ is 0.010.010.01, and the computed advantages are {$,1.0\}$ 1.0 }, which overly magnifies the minor numerical fluctuation. To mitigate this issue, we propose to clip $\mathrm{std}(R)$ to be at least $\epsilon_{std}$ to ensure that the denominator does not become too small:
@@ 4.1 Experimental Setup
[…]
@@ 5.2 RAG with Context Compression
RAG systems usually concatenate all retrieved passages as the context of LLMs. However, this may introduce data noise and computational overhead due to imperfect retrieval and overlong context. Recently, many works have attempted to compress context and retain important information, mainly including two categories: (1) Reranking methods, which rerank retrieved passages and retain top-ranked passages Hwang et al. (2024b); Mortaheb et al. (2025); Mao et al. (2021); Chen et al. (2025) (2) Summarization methods, which extract relevant information from retrieved passages and consolidate them into a narrative Wang et al. (2023); Zhao et al. (2024a); Zhu et al. (2024); Xu et al. (2024). While effective, existing works neglect evidence reasoning, which risks filtering out key clues and struggles with generalization.